### v1

# Scaling LLM Planning: NL2FLOW for Parametric Problem Generation and Rigorous Evaluation

Abstract
Progress in enhancing large language model (LLM) planning and reasoning capabilities is significantly hampered by the bottleneck of scalable, reliable data generation and evaluation. To overcome this, I introduce NL2FLOW, a fully automated system for parametrically generating planning problems – expressed in natural language, a structured intermediate representation, and formal PDDL – and rigorously evaluating the quality of generated plans. I demonstrate NL2FLOW’s capabilities by generating a dataset of 2296 problems in the automated workflow generation domain and evaluating multiple open-sourced, instruct-tuned LLMs. My results reveal that the highest performing models achieved 86% success in generating valid plans and 69% in generating optimal plans, specifically for problems with feasible solutions. Regression analysis shows that the influence of problem characteristics on plan generation is contingent on both model and prompt design. Notably, I observed that the highest success rate for translating natural language into a JSON representation of a plan was lower than the highest rate of generating a valid plan directly. This suggests that unnecessarily decomposing the reasoning task – introducing intermediate translation steps – may actually degrade performance, implying a benefit to models capable of reasoning directly from natural language to action. As I scale LLM reasoning to increasingly complex problems, the bottlenecks and sources of error within these systems will inevitably shift. Therefore, a dynamic understanding of these limitations – and the tools to systematically reveal them – will be crucial for unlocking the full potential of LLMs as intelligent problem solvers.
1 Introduction
Figure 1: Steps to obtain a plan from a planning problem in natural language
Planning is crucial for modern conversational AI systems, enabling them to achieve user goals by leveraging available resources such as APIs, agents, tools, databases, and internet. While complex planning systems often rely on symbolic planners, building such systems is a significant undertaking. A key challenge lies in knowledge engineering – translating natural language problem descriptions into formal representations suitable for planning algorithms. This translation process is prone to error [1] and can become a bottleneck as system complexity increases. Furthermore, classical planning approaches, based on algorithms like those implemented in Planning Domain Definition Language (PDDL), can struggle with scalability.
Classical AI planning, while offering guarantees like optimality, requires precise formalization of problems. Errors in translating natural language into PDDL can lead to planning failures. Large Language Models (LLMs) offer a promising alternative, demonstrating capabilities in language understanding, generation, and reasoning. However, LLM-based planning remains unreliable, and current approaches often focus on free-form plan generation or reinforcement learning within limited environments [2]. A significant advancement would be to leverage LLMs to reliably translate natural language planning problems into formal representations, or directly generate plans with high logical consistency (Figure 1). If an LLM can accurately translate or generate plans, the complexity of building a robust natural language planning system is drastically reduced.
However, progress in improving LLM planning and translation capabilities is hindered by a lack of scalable data generation and evaluation methods. To address this, I introduce NL2FLOW, a system designed to 1) generate planning problems expressed in natural language, an intermediate representation, and PDDL, and 2) evaluate the quality of generated plans. This study leverages NL2FLOW to investigate the inherent planning capabilities of open-source LLMs, establishing a baseline performance without task-specific optimization or architectural modifications. My primary goal is to identify domains where LLMs exhibit natural planning strengths and to develop a data generator capable of producing scalable, configurable planning datasets relevant to real-world applications. This will allow researchers to systematically probe the limitations of LLMs in complex planning scenarios, drawing on insights from automated planning to identify structural weaknesses in LLM-based approaches. Ultimately, this work aims to inform the development of improved architectures and training techniques for deep neural networks designed for planning tasks.
In this paper, I focus on applying LLMs to simple planning problems within the domain of workflow generation. Workflow creation – decomposing complex problems into structured, sequential tasks – is a common daily activity. I define “simple” problems as those solvable by symbolic planners within a short timeframe (e.g., a few seconds). I explore two primary approaches: 1) using LLMs to translate natural language problem descriptions into structured formats (e.g., JSON) suitable for symbolic planners, and 2) examining the direct generation of plans by LLMs from natural language descriptions.
This study is impactful in several key aspects:
  • Scalable Data Generation: NL2FLOW generates data randomly and parametrically, enabling scalable dataset creation. Scalability is primarily limited by the computational capacity of the symbolic planner used for data validation, ensuring the generated data is symbolically verifiable. The parametric nature of data generation also provides well-defined explanatory variables for downstream analysis. Critically, this approach requires no human or LLM intervention in data creation or inspection, making it ideal for training deep neural networks that require large datasets and programmatic reward computation.
  • Multi-Format Data: NL2FLOW data includes natural language descriptions, enabling the training and evaluation of LLMs for planning tasks, intermediate representations in python object, and PDDLs.
  • Symbolic Evaluation: The inclusion of PDDL allows for the symbolic evaluation of LLM-generated plans, providing mathematically guaranteed metrics like optimality. This enables scalable and reliable evaluation.
  • Problem Complexity Metric: I define problem complexity as the length of the optimal plan, simplifying downstream analysis and providing a clear measure of difficulty.
2 Related Work
Current evaluation of reasoning capacity often focuses on final answer accuracy for benchmarks like MATH500 [3], AIME24, GSM8K [4], and MATH. However, this metric overlooks the validity of the reasoning ’process’ itself; a correct answer can be achieved through flawed reasoning, effectively rewarding a successful, yet invalid, plan. Furthermore, relying solely on final answer correctness leaves the quality and structure of the reasoning process opaque. Similarly, BIRD [5], a natural language to SQL translation benchmark, primarily relies on final answer accuracy, despite the inherent planning aspects of SQL query construction alongside elements like intent detection. GSM-Symbolic [6] attempts to address this by generating mathematical problems from symbolic templates, allowing for manual verification of solutions. ReasonAgain [7] leverages GPT-4o to generate Python programs for GSM8K and MATH, validating reasoning by executing the generated code.
Beyond these benchmarks, several studies utilize “puzzle” datasets [8] - Tower of Hanoi, river crossing, blocks world, and checker jumping - to assess LLM reasoning. These puzzles offer the advantage of verifiable reasoning steps through simulation and can be scaled to challenge LLMs as they saturate. However, the applicability of these abstract puzzles to real-world scenarios remains limited.
A growing body of work focuses on evaluating and training Large Language Models (LLMs) for workflow automation and reasoning. Several datasets have been proposed to assess LLM planning capabilities, differing in their generation methods (static vs. dynamically generated) and evaluation approaches (manual vs. programmatic). This section details relevant work, highlighting these key distinctions.
2.1 Static Datasets & Primarily Manual/LLM-as-Judge Evaluation
FlowBench [9] presents a static dataset of 51 scenarios across 6 domains, focusing on multi-turn interactions. Evaluation relied heavily on human verification by comparing expected and observed traces, identifying a need for improved LLM planning. Similarly, HuggingGPT [10] utilizes a static, human-annotated dataset augmented with GPT-4 generated pseudo-labels, and employs LLM-as-a-judge for graph tasks. TaskLAMA [11] also provides a static dataset, created through human annotation, for structured complex task decomposition. ISG-BENCH [12] focuses on interleaved text and image generation, utilizing a static dataset and relying on MLLM-as-a-judge for scoring alignment, coherence, and helpfulness, grounded in human-annotated truth. WikiPlan and RecipePlan contribute static text-image pairs for task completion. These datasets, while valuable, often require significant human effort for creation and/or evaluation [13].
2.2 Static Datasets & Programmatic Evaluation
TaskBench [14] divides task automation into decomposition, tool selection, and parameter prediction, leveraging a static dataset of 17,331 samples generated with GPT-4. The authors employed programmatic metrics (F1, Normalized Edit Distance, Rouge) for evaluation, with some partial manual inspection. This approach offers scalability due to a large sample size but can be limited by the quality and coverage of the initial dataset. Open Grounded Planning [2] also aims to generate executable plans, providing LLMs with an extensive action library and leveraging existing datasets. Evaluation combines programmatic metrics (executability, pass rate) with LLM-as-a-judge for assessing completeness, feasibility, and relevance.
2.3 Algorithmically Generated Datasets & Programmatic Evaluation
WORFBENCH [15] represents a step towards automated evaluation through its data generator for graph workflow structures. LLM responses are evaluated via subgraph matching. However, there is a limitation: the need to define the entire graph upfront, hindering scalability – a limitation that search-based methods, like my own, avoid by not requiring complete graph definition for data generation and evaluation.
My work builds upon these efforts by presenting a fully automated pipeline for both data generation and evaluation, addressing the limitations of static datasets and the need for scalable, programmatic (symbolic) assessment of LLM reasoning and planning capabilities.
3 Method
This section details the automated data generation and plan evaluation pipeline used in this study. I leveraged NL2FLOW (https://github.com/IBM/nl2flow) to generate a dataset of planning problems and evaluate plans produced by large language models.
3.1 Domain Description
This work explores automated planning within the domain of business process automation, specifically focusing on task flow generation. The system’s objective is to autonomously determine and execute a sequence of actions to achieve user-defined goals, such as submitting an expense report. I model each action as a function requiring a set of input parameters. Initially, the values of these parameters are often unknown, representing information the system needs to acquire. To address this, the system employs a dialogue-based approach. When a parameter value is missing, the system initiates a request for information via an ’ask’ action – a dialogue act designed to elicit the necessary value from the user. This process of identifying and requesting missing parameter values is known as slot-filling, a common technique in task-oriented dialogue systems. Successfully executing an action updates the system’s internal state, transitioning the corresponding parameter from an unknown to a known state. This allows the system to proceed with subsequent actions in the plan.
3.2 Planning Data Generation
Table 1: Planning data generator configuration
Parameters Values
Actions [2, 3, 4, 5]
Arity [1, 2, 3]
Coupling [0, 0.375, 0.626, 1.0]
Slot-fillable variables [0.375, 0.626, 1.0]
Goals [1, 2]
Figure 2: Distribution of the number of steps in optimal plans across the dataset
I generated 2,296 planning data (intermediate representation + natural language descriptions + PDDL) using a parameterized approach (Table 1) to control the complexity and characteristics of the generated problems. Key parameters included the number of actions, action arity (the number of input and output parameters per action), the number of goal actions, the proportion of slot-fillable variables (variables requiring values to be determined during planning), and the proportion of coupled actions (actions where the output of one action serves as input to another). 30 variables were generated and designated as either slot-fillable or initially known. Actions were then created, establishing dependencies between them by connecting action outputs to subsequent action inputs. Variables not directly used within any action were assigned to the available data, whose values are already known. An action is executable only when all required input parameter values are known; upon execution, the values of its output parameters become known. Goal actions were randomly selected from the generated action set. The resulting intermediate representation (python object) was compiled into PDDL strings using NL2FLOW, facilitating symbolic verification of large language model plans. Each generated problem instance was assigned a unique hash for identification and deduplication. The optimal plan length for the generated planning data ranged from 0 to 8 (Figure 2). A plan consisting of only one action was not feasible. Because the minimum arity of any action in this domain was 1, each action required at least one input variable. This input variable must be obtained either by invoking the ‘ask‘ action or another action, precluding a valid plan of length 1.
3.3 Prompt Generation
Listing 1: Verbose planning prompt
CONTEXT
The system has actions a__0, a__1, and a__2.
Action a__1 has input parameters v__13, v__12, and v__11 and it outputs variables v__15, v__16, and v__18. To execute action a__1, variables v__13, v__12, and v__11 must be known. After executing action a__1, variables v__15, v__16, and v__18 are known.
Action a__2 has
INSTRUCTION
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as no plan
Here are some examples.
Example #1
Now output your plan explanation and plan.
{plan_explanation_and_plan}
Listing 2: Concise planning prompt
CONTEXT
The system has actions a__0, a__1, and a__2. To execute an action the values of their inputs must be known.
actions:
name: a__1
inputs: v__11, v__12, v__13
outputs: v__15, v__18, v__16
name: a__2
inputs:
INSTRUCTION
Develop a concise plan to achieve the goals. The shorter plan is preferable. A plan consists of individual actions, each represented by its outputs and an action name followed by its parameters in parentheses (e.g., outputs = <action-name>(<parameters>). When it becomes impossible to achieve the objectives, the appropriate course of action should be considered as no plan
Here are some examples.
Example #1
Now output your plan explanation and plan.
{plan_explanation_and_plan}
Listing 3: Planning response example
PLAN EXPLANATION
0. Execute action a__2. This will result in acquiring v__1. aa is required later by action A.
1. Execute action a__1 with aa as input. This will result in acquiring v__2. v__2 is required later by action a__3. Since executing A was a goal of this plan, return the results of A(v__1) to the user.
2. Execute action a__3 with bb as input. Since executing a__3 was a goal of this plan, return the results of a__3(v__2) to the user.
PLAN
[0] v__1 = a__2()
[1] v__2 = a__1(v__1)
[2] a__3(v__2)
Listing 4: Translation prompt
CONTEXT
The system has actions a__0, a__1, and a__2.
Action a__1 has input parameters v__13, v__12, and v__11 and it outputs variables v__15, v__16, and v__18. To execute action a__1, variables v__13, v__12, and v__11 must be known. After executing action a__1, variables v__15, v__16, and v__18 are known.
Action a__2 has
INSTRUCTION
Write a json to describe the context.
Here are some examples.
Example #1
Now output your json.
{json}
This study examined the capacity of large language models (LLMs) to translate natural language problem descriptions into JSON schemas and to generate executable plans to achieve specified goals. I adopted a few-shot learning approach, constructing prompt templates for both tasks and providing six example demonstrations. For the planning task, I investigated the impact of prompt length, specifically comparing ’verbose’ and ’desclarative’ prompt templates. The ’verbose’ prompts (mean length 13,413 characters) included a detailed, natural language description of all relevant elements within the planning problem, aiming to provide the LLM with comprehensive contextual information. Conversely, ’concise’ prompts (mean length 9,674 characters) presented these elements as a concise, enumerated list, minimizing verbose text and focusing on key parameters. This variation was designed to assess the LLM’s sensitivity to the level of contextual detail provided in the prompt. For translation, a single prompt template (see Listing 4) was used, designed to elicit accurate and fluent translations. The complete prompt templates used for each task – verbose_planning_prompt.txt, concise_planning_prompt.txt, and translation_prompt.txt – are included in the supplementary materials to facilitate reproducibility, enable independent verification of my results, and encourage further research in this area.
Each prompt template followed a consistent structure comprising three sections: Context, Instruction, and Examples (Listings 1 and 2). The Context section provided a programmatically generated natural language description of the problem. The Instruction section detailed the desired planning process, specifying preferences for plan structure, the required output format, and guidance for handling unsolvable problems. To facilitate parsing and execution, plans were to be expressed in a Python-like syntax (e.g., outputs = ⟨action name⟩(⟨parameters⟩)). I standardized action names as a__⟨number⟩and variable names as v__⟨number⟩to promote consistency and reduce ambiguity.
Crucially, each example within the prompt presented a natural language description of the plan execution before the corresponding plan expressed in the target format (Listing 3). This was intended to guide the LLM by demonstrating the logical connection between action sequences and their executable representations, thereby increasing the likelihood of generating valid and goal-directed plans. Example prompts are provided in the supplementary materials.
3.4 Language Model Response Parsing
To establish a lower bound on large language model performance, I employed minimal parsing techniques for both translation and planning tasks, prioritizing simplicity and avoiding potentially biasing post-processing steps. For the translation task, model outputs were structured as JSON strings representing a planning problem. These strings were then parsed using PyDantic, a data validation and settings management library, to ensure adherence to a predefined schema. This structured format enabled automated evaluation metrics and facilitated downstream analysis. For the planning task, each action within a generated plan was extracted and stored as a string within a Python list. This list was subsequently used as input to the NL2FLOW debugger, a symbolic planning validation tool that assesses the logical consistency and feasibility of the generated plan.
3.5 Plan Validation
To assess the soundness of generated plans, the pddl compilation omits goal achievement. My validation approach distinguishes between soundness and validity. For soundness checking, I compiled the planning problem into PDDL, but instead of including the original goals as requirements, I incorporated the actions reported as executed within the LLM’s response. This approach tests the LLM’s claimed actions actually executable. Validity is checked by compiling with both original goals and actions in LLM’s plan as requirements; failure to achieve goals indicates a non-valid plan where a planner finds a longer optimal plan.
To quantify plan optimality, I compare the length of the LLM-generated plan to the length of optimal plans derived using the $\ast$ planner. I formulate the planning problem in PDDL using only the original goals as requirements and leverage $\ast$ to obtain a set of optimal plans. Following a successful validity check, an LLM-generated plan is considered sub-optimal if its length exceeds the length of the shortest $\ast$-derived optimal plan. This comparison provides a quantitative measure of how efficiently the LLM solves the planning problem relative to a state-of-the-art planner.
I investigated the use of VAL [16] for plan validation; however, VAL does not currently support the PDDL language features necessary to represent the complexities introduced by my NL2FLOW compilation process, precluding its use in this evaluation.
4 Results
4.1 Planning Task
4.1.1 Plan Length
[…]
Figure 3: The distribution of plan lengths (number of steps) generated by LLMs
4.1.2 Sound, Valid, and Optimal Planning
Overall, Llama-3.3-70B-instruct, when paired with the verbose prompt template, demonstrated the strongest planning performance (Figure 4). This model achieved the highest rates of generating sound plans (80%), valid plans (77%), and optimal plans (64%) among all tested models. Utilizing the concise prompt template resulted in a noticeable decrease in performance, with rates dropping to 71%, 60%, and 46% respectively. A similar, though more pronounced, trend was observed for Llama-3.1-8B-instruct. For clarity, a sound plan is defined as one that is executable; a valid plan is executable and achieves the goal state; and an optimal plan is executable, reaches the goal state, and does so with minimal cost.
4.1.3 No Plan
A key limitation observed across all models was the difficulty in correctly identifying unsolvable problems (Figure 5) – those for which no plan exists. The superior performance of Llama-3.3-70B-instruct in generating sound, valid, and optimal plans was largely attributable to its enhanced ability to recognize these ‘no plan’ scenarios. Analysis of explanatory variables – including the number of actions, arity, coupling, the proportion of slot-fillable variables, and the number of goals – revealed differing factors contributing to failures in identifying ‘no plan’ problems depending on the prompt template (Tables LABEL:tab:lrnoplanlong and LABEL:tab:lrnoplanshort). With the verbose prompt template, arity was the sole consistent predictor of failure (higher arity correlated with a higher rate of incorrectly attempting to generate a plan when none existed). Conversely, using the concise prompt template, the number of goals was the primary factor, with a higher number of goals increasing the likelihood of failing to identify ‘no plan’ problems.
Figure 4: Comparison of LLM performance in generating sound, valid, and optimal plans using verbose and concise prompt templates
Figure 5: Comparison of large language model performance (accuracy) in detecting ’no plan’ scenarios, evaluated with verbose and concise prompt templates
4.1.4 Feasible Plan
[…]
Regarding optimal plan generation, coupling consistently explained a significant portion of the variance across all models when the verbose prompt was used (Table LABEL:tab:lrplanoptimallong). This trend was observed only in Llama-3.3-70B-instruct when using the concise prompt (Table LABEL:tab:lrplanoptimalshort). These results suggest that the impact of problem characteristics on plan generation is both model- and prompt-dependent.
Figure 6: Performance comparison of large language models on planning tasks with solutions
4.2 Natural Language to JSON Translation
Llama-3.1-405B-Instruct-FP8 achieved the highest success rate in generating valid JSON strings (99%) and minimizing errors in content (Figure 7). Specifically, 73% of the JSON outputs from this model contained all required elements without hallucinated information. Codellama-34B-Instruct-HR and Llama-3.3-70B-Instruct consistently failed to generate valid JSON. Across all models, the primary source of translation error was the omission of required elements, rather than the inclusion of hallucinated information. Linear regression analysis revealed a negative correlation between the proportion of correctly translated elements and both the number of actions and their arity (Table LABEL:tab:olstranslation). Conversely, a positive correlation was observed between correct translation and both the degree of coupling between actions and the proportion of slot-fillable variables.
Figure 7: LLM-based natural language to JSON translation for planning
5 Conclusion
This study introduced a fully automated pipeline for generating and evaluating reasoning data, leveraging a planning data generator capable of producing both PDDL and intermediate Python object representations alongside a symbolic evaluator for assessing plan validity and JSON translation accuracy. This approach enables scalable data creation and rigorous assessment crucial for advancing large language models with reasoning and planning capabilities.
My evaluation, using simple planning problems generated by NL2FLOW, revealed that current open-sourced, instruct-tuned LLMs demonstrate a surprising capacity for direct plan generation. Notably, I observed that the highest success rate for translating natural language into a JSON representation of a plan was lower than the highest rate of generating a valid plan directly. This suggests that unnecessarily decomposing the reasoning task – breaking it down into intermediate translation steps – may actually hinder performance, implying a benefit to models learning to reason directly from natural language to action. Ultimately, the bottlenecks and sources of error within a reasoning system are likely to shift depending on the complexity and scale of the problem. As I push the boundaries of LLM reasoning, understanding these dynamic limitations will be paramount.
Further analysis through regression revealed that the impact of problem characteristics on plan generation is significantly dependent on the specific language model and the prompting strategy employed. This highlights the critical need for model-specific insights; what works for one LLM may not generalize to another.
6 Acknowledgements
The author developed NL2FLOW in close collaboration with Dr. Tathagata Chakraborti. Dr. Chakraborti provided invaluable guidance throughout the design and implementation of the system, particularly in shaping the core planning algorithm. The author also gratefully acknowledges Dr. Junkyu Lee at IBM Research for providing critical feedback on an early draft of this manuscript.